OpenStudy (anonymous):

Margie calculated that she would spend $175 on school supplies this year. She actually spent $97.50 on school supplies. What is Margie’s percent of error? 77.3% 79.5% 44.2% 4.4%

OpenStudy (anonymous):

@johnweldon1993 !!!!!!

OpenStudy (anonymous):

I want to make sure im right!

OpenStudy (johnweldon1993):

What'd you get? :)

OpenStudy (anonymous):

B??

OpenStudy (anonymous):

79.5?

OpenStudy (johnweldon1993):

How did you get that?

OpenStudy (johnweldon1993):

Ehh nvm, I just like showing the process haha you're correct :P

OpenStudy (anonymous):

\[\frac{ 77.5 }{ 97.5 }\times100 =79.48717949 \]

OpenStudy (johnweldon1993):

\[\large \color \red \checkmark\] BAM!

OpenStudy (richgall____):

its actually C

OpenStudy (richgall____):

i was looking at similar problems like yours and most ppl cam up with 44.2

OpenStudy (anonymous):

but also \[175-97.50=77.5\]

OpenStudy (johnweldon1993):

\[\large \text{% Error} = \frac{\text{Theoretical Value - Experimental Value}}{\text{Theoretical Value}}\]

OpenStudy (richgall____):

yea ik but u have to find the percent of what 77.5 was of that

OpenStudy (richgall____):

why would they make the problem that simple i mean come on

OpenStudy (anonymous):

\[\frac{ 77.5 }{ 97.5 }=79.48717949\] there you go :P :)

OpenStudy (richgall____):

now i understand it better lol okay ;)

OpenStudy (johnweldon1993):

I believe what other people are thinking...is \[\large \frac{97.5}{175} = 55.7 \text{ Percent}\] That is the amount she was correct...so they think 1 - that = 44.3 would be the percent she is wrong...but that is an incorrect observation

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

yup :) <3